# How (Mis)calibrated is Your Federated CLIP and What To Do About It?

### v1

@@ A.4 Calibration baselines
Eq. 8 computes an auxiliary calibration loss by taking, over $m$ samples in a minibatch, that penalizes the mismatch between observed mean correctness and mean confidence, thereby guiding the model to better align its confidence estimates with empirical accuracy, without strict constraint on each example. Note that during backpropagation, the gradients are backpropagated only through the confidence term $s_{i}$.
@@
Additional implementation details. In this paragraph, we elaborate on the implementation details, providing additional information beyond what is included in the main paper. In $\text{FL}^{2}\text{oRA}$, we only fine-tune the LoRA-adapted layers of the transformer blocks, while keeping all original weights frozen. Specifically, for both the image and text encoders, only the LoRA matrices $W_{A}$ and $W_{B}$ of the query, key, and value projections in the attention modules are trainable. All other parameters, including projection, MLP, and layer normalization weights, remain frozen to reduce memory footprint and improve training efficiency. We keep the rank of LoRA metrices to 2 and a dropout rate of 0.25 to regularize the input of the LoRA modules.
Appendix B Additional empirical evaluation
In this section we report additional experimental analyses and detailed breakdown of the results presented in the main paper.
B.1 Effect of Temperature Scaling
In Figure 7, we show how the temperature hyperparameter $\tau$ effects the model calibration (ECE) on CIFAR-10 and CIFAR-100 datasets using the post-hoc calibration technique temperature scaling [guo2017calibration]. The results showcase moderate temperatures yield better calibration, while too small or too large $\tau$ harms it. ZS-CLIP is highly sensitive to $\tau$ as ECE improves markedly near $\tau=0.5$ but degrades rapidly as $\tau$ increases, indicating that mild softening corrects overconfidence whereas excessive softening induces underconfident predictions. Among federated baselines, FedOTP and FedPGP tend to worsen as $\tau$ grows, suggesting a monotonic loss of calibration with over-smoothing, whereas PromptFL and FedPHA are comparatively stable with smaller ECE variation. However, our proposed $\text{FL}^{2}\text{oRA}$ maintains both the lowest and most stable ECE over a wide range of $\tau$, pointing to robustness to the temperature choice and reduced reliance on per-dataset tuning.
(a) CIFAR-10
(b) CIFAR-100
Table 8: Comparison of federated CLIP fine-tuning methods. Results have been reported on the in-distribution setting using datasets: CIFAR-10 and CIFAR-100 in terms of Brier score [brierscore] (Mean squared probability error), NLL [nll] (Negative LogLikelihood).
CIFAR-10 CIFAR-100
Method Brier score $\downarrow$ NLL $\downarrow$ Brier Score $\downarrow$ NLL $\downarrow$
ZS-CLIP [radford2019language] 0.17 0.37 0.48 1.40
FL setting (without calibration)
PromptFL [promptfl] 0.12 0.23 0.36 0.93
FedPGP [fedpgp] 0.13 0.28 0.42 1.12
FedOTP [fedotp] 0.09 0.20 0.38 1.00
FedPHA [fedpha] 0.10 0.20 0.35 0.91
$\text{FL}^{2}\text{oRA}$-text (Ours) 0.11 0.23 0.36 0.90
$\text{FL}^{2}\text{oRA}$-vision (Ours) 0.05 0.09 0.24 0.57
$\text{FL}^{2}\text{oRA}$ (Ours) 0.05 0.10 0.50 1.43
FL setting (with calibration: DCA [dca])
PromptFL [promptfl] 0.12 0.24 0.36 0.92
FedPGP [fedpgp] 0.13 0.28 0.38 1.00
FedOTP [fedotp] 0.09 0.20 0.37 0.97
FedPHA [fedpha] 0.10 0.20 0.34 0.90
$\text{FL}^{2}\text{oRA}$-text (Ours) 0.11 0.23 0.36 0.90
$\text{FL}^{2}\text{oRA}$-vision (Ours) 0.05 0.09 0.24 0.57
$\text{FL}^{2}\text{oRA}$ (Ours) 0.05 0.09 0.23 0.52
FL setting (with calibration: MDCA [mdca])
PromptFL [promptfl] 0.12 0.23 0.36 0.92
FedPGP [fedpgp] 0.14 0.30 0.38 0.98
FedOTP [fedotp] 0.09 0.20 0.38 1.00
FedPHA [fedpha] 0.10 0.20 0.34 0.91
$\text{FL}^{2}\text{oRA}$-text (Ours) 0.11 0.23 0.36 0.90
$\text{FL}^{2}\text{oRA}$-vision (Ours) 0.05 0.09 0.24 0.57
$\text{FL}^{2}\text{oRA}$ (Ours) 0.05 0.09 0.23 0.52
B.2 Additional calibration metrics
@@
In Table 8, we present the detailed Brier score and NLL calibration results for the various $\text{FL}^{2}\text{oRA}$ variants as well as the baseline methods. We observe that both $\text{FL}^{2}\text{oRA}$ and $\text{FL}^{2}\text{oRA}$-vision consistently achieve better calibration, while $\text{FL}^{2}\text{oRA}$-text performs comparably to the baselines. This is consistent with the findings in the main paper, where for all the calibration metrics, $\text{FL}^{2}\text{oRA}$ and it variants were better and/or competitive without needing extra regularization losses.
@@
(a) CIFAR-10
(b) CIFAR-100
Figure 9: Analysis of $|w_{\text{zs}}-w_{\text{LoRA}}|$ over each communication round. We report the average results of all LoRA trainable encoder layers of each clients on (a) CIFAR-10 and (b) CIFAR-100 datasets.
B.3 Impact of client participation rate
@@
Figure 10: Comparison of number of trainable parameters. We show the qualitative results of Accuracy vs. ECE metrics on CIFAR-10 and CIFAR-100 datasets, where the markers’ size define the number of trainable parameters of each method.
B.4 Additional insights on improved calibration
In Section 4.4 of the main paper, we hypothesize that $\text{FL}^{2}\text{oRA}$ calibration gains stem from LoRA’s constrained weight updates: by preventing large shifts in weight norms and logit scales across clients, LoRA implicitly enhances the ensemble effect induced by FedAvg. Figure 9 empirically supports this hypothesis on CIFAR-10/100. Specifically, for each client, we measure the weight drift relative to the pretrained CLIP model before each communication round, averaging the drift over all LoRA layers. As shown in the figure, the inter-client standard deviation is negligible, indicating that LoRA indeed limits parameter drift and keeps logit scales stable.
B.5 Complexity analysis
In Figure 10, we qualitatively illustrate the trade-off between trainable parameters, accuracy, and ECE calibration on CIFAR-10 and CIFAR-100, both without and with DCA. Here, the trainable parameters are defined by markers and the marker size depends on the number of the training parameters. The FL prompt tuning methods with moderate trainable parameters showcase substantial accuracy gains and noticeably lower ECE, showing strong returns for minimal complexity. However, the LoRA variants in our proposed $\text{FL}^{2}\text{oRA}$ push the frontier further. As the parameter budget increases (from text to vision and both), it moves toward higher accuracy and lower ECE, though improvements become incremental, indicating diminishing returns at the high-compute end. However, applying DCA on them showcase similar calibration. Finally, $\text{FL}^{2}\text{oRA}$ comes up with a trade-off between training parameter size and performance, but the improved calibration showcase a positive indication in the FL settings.
B.6 Detailed results of the main tables
In Table 9, 10 and 11, we provide the detailed results of individual datasets PACS, OfficeHome and VLCS respectively in the domain-generalization setting. The average results of them are reported in the Table. 4 of the main paper. Similarly in Table 12, 13, 14, 15 and 16, we provide the detailed results of Food101, DTD, Caltech101, Flowers102, and OxfordPets datasets, corresponding to Table 5 of the main paper, for the base-to-new generalization setting.
@@ B.6 Detailed results of the main tables
$\text{FL}^{2}\text{oRA}$-text (Ours) 98.78 0.96 0.37 0.76 99.29 1.35 0.46 0.95 99.85 0.92 0.40 0.68
$\text{FL}^{2}\text{oRA}$-vision (Ours) 99.40 0.81 0.34 0.61 99.82 0.66 0.32 0.40 99.90 0.39 0.17 0.26
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 98.65 0.75 0.22 0.50 99.32 1.09 0.38 0.44 99.90 0.54 0.18 0.33
$\text{FL}^{2}\text{oRA}$-vision (Ours) 99.41 0.64 0.29 0.52 99.82 0.60 0.28 0.31 99.90 0.33 0.18 0.21
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 98.78 0.90 0.26 0.42 99.24 0.89 0.32 0.58 99.85 0.74 0.29 0.41
$\text{FL}^{2}\text{oRA}$-vision (Ours) 99.41 0.64 0.31 0.55 99.82 0.64 0.30 0.36 99.90 0.37 0.20 0.24
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 93.22 1.64 0.39 0.81 97.78 1.22 0.40 0.80
$\text{FL}^{2}\text{oRA}$-vision (Ours) 98.18 0.77 0.31 0.58 99.32 0.66 0.29 0.46
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 94.10 1.18 0.32 0.91 97.99 0.89 0.28 0.55
$\text{FL}^{2}\text{oRA}$-vision (Ours) 98.03 0.78 0.32 0.59 99.29 0.59 0.27 0.41
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 93.65 1.57 0.48 0.68 97.88 1.03 0.34 0.52
$\text{FL}^{2}\text{oRA}$-vision (Ours) 98.23 0.68 0.31 0.54 99.34 0.58 0.28 0.42
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 87.52 3.68 0.83 3.00 75.65 4.09 0.92 4.50 94.94 2.89 0.77 2.16
$\text{FL}^{2}\text{oRA}$-vision (Ours) 87.94 3.27 1.00 2.27 81.15 2.89 1.18 2.69 94.21 1.51 0.69 1.60
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 88.36 3.64 0.91 2.92 76.14 3.59 1.06 3.02 95.17 2.16 0.66 1.87
$\text{FL}^{2}\text{oRA}$-vision (Ours) 88.49 3.97 1.24 3.33 80.97 4.55 1.35 3.53 94.32 2.16 0.76 1.83
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 87.52 3.50 0.93 3.02 75.55 4.32 0.92 4.00 94.94 2.58 0.75 2.10
$\text{FL}^{2}\text{oRA}$-vision (Ours) 87.94 3.51 1.11 2.16 80.92 3.33 1.17 2.29 94.21 1.54 0.55 1.61
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 93.45 2.96 0.87 2.56 87.89 3.41 0.85 3.06
$\text{FL}^{2}\text{oRA}$-vision (Ours) 92.96 2.37 0.85 2.17 89.06 2.51 0.93 2.18
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 93.38 2.72 0.81 2.30 88.26 3.03 0.86 2.53
$\text{FL}^{2}\text{oRA}$-vision (Ours) 92.97 2.23 0.52 1.92 89.19 3.23 0.97 2.65
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 93.56 2.75 0.84 2.64 87.89 3.29 0.86 2.94
$\text{FL}^{2}\text{oRA}$-vision (Ours) 93.08 2.30 0.71 2.26 89.04 2.67 0.89 2.08
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 100.00 2.64 2.32 2.64 80.40 4.47 1.23 5.00 90.41 3.96 1.38 3.62
$\text{FL}^{2}\text{oRA}$-vision (Ours) 99.64 1.80 0.97 1.50 83.27 7.00 2.38 6.63 90.99 2.31 0.64 2.47
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 100.00 2.06 1.86 2.05 77.83 5.56 2.24 7.24 90.43 3.59 1.24 3.06
$\text{FL}^{2}\text{oRA}$-vision (Ours) 99.64 1.74 0.85 1.33 82.70 7.71 2.58 6.71 90.86 2.59 0.83 2.10
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 100.00 1.36 1.26 1.31 79.03 5.55 2.24 6.29 91.18 2.57 0.82 2.69
$\text{FL}^{2}\text{oRA}$-vision (Ours) 99.64 1.74 0.94 1.43 82.29 5.73 2.28 6.26 91.43 2.02 0.62 2.02
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 84.43 4.46 1.48 4.33 88.81 3.88 1.60 3.90
$\text{FL}^{2}\text{oRA}$-vision (Ours) 87.74 2.92 1.18 2.59 90.41 3.51 1.29 3.30
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 83.58 4.71 1.72 5.02 87.96 3.98 1.77 4.34
$\text{FL}^{2}\text{oRA}$-vision (Ours) 87.52 3.39 1.08 2.16 90.18 3.86 1.34 3.08
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 84.19 3.26 1.03 3.99 88.60 3.18 1.34 3.57
$\text{FL}^{2}\text{oRA}$-vision (Ours) 88.22 2.65 1.11 3.20 90.40 3.04 1.24 3.23
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 90.06 91.23 90.64 0.49 0.56 0.52 0.15 0.19 0.17 0.37 0.52 0.43
$\text{FL}^{2}\text{oRA}$-vision (Ours) 89.29 90.58 89.93 1.14 0.74 0.90 0.30 0.22 0.25 1.21 0.67 0.86
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 90.13 91.24 90.68 0.66 0.46 0.54 0.18 0.12 0.14 0.64 0.36 0.46
$\text{FL}^{2}\text{oRA}$-vision (Ours) 89.22 90.54 89.87 1.75 0.55 0.84 0.36 0.15 0.21 1.73 0.53 0.81
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 90.08 91.28 90.68 0.51 0.47 0.49 0.14 0.17 0.15 0.36 0.49 0.42
$\text{FL}^{2}\text{oRA}$-vision (Ours) 89.26 90.52 89.89 1.23 0.72 0.91 0.33 0.20 0.25 1.26 0.67 0.87
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 79.51 58.49 67.40 5.36 16.78 8.12 1.62 4.80 2.42 4.91 16.25 7.54
$\text{FL}^{2}\text{oRA}$-vision (Ours) 81.34 48.59 60.84 6.97 17.14 9.91 1.81 3.89 2.47 7.45 16.95 10.35
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 79.27 58.00 66.99 6.47 16.98 9.37 2.05 5.37 2.97 5.98 16.06 8.72
$\text{FL}^{2}\text{oRA}$-vision (Ours) 81.92 48.50 60.93 3.81 19.14 6.35 1.16 4.38 1.83 4.83 18.81 7.68
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 79.78 55.62 65.54 4.91 17.38 7.66 1.50 3.97 2.18 4.61 17.03 7.26
$\text{FL}^{2}\text{oRA}$-vision (Ours) 81.60 48.66 60.96 6.63 17.27 9.58 1.79 3.95 2.46 7.24 17.07 10.17
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 98.17 93.32 95.68 1.62 2.92 2.08 0.58 1.07 0.75 1.17 2.44 1.58
$\text{FL}^{2}\text{oRA}$-vision (Ours) 98.40 92.78 95.51 2.91 7.21 4.15 0.91 3.57 1.46 2.62 4.09 3.20
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 98.17 93.61 95.84 1.33 2.80 1.80 0.49 0.92 0.64 0.91 2.18 1.28
$\text{FL}^{2}\text{oRA}$-vision (Ours) 98.41 92.82 95.53 2.32 6.57 3.43 0.84 3.75 1.37 1.95 4.15 2.65
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 98.17 93.45 95.75 1.62 2.65 2.01 0.57 0.90 0.70 1.13 2.20 1.49
$\text{FL}^{2}\text{oRA}$-vision (Ours) 98.36 92.79 95.50 2.76 6.18 3.82 0.90 3.59 1.44 2.50 4.12 3.11
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 97.63 72.67 83.32 3.51 9.02 5.05 1.25 1.95 1.52 3.26 9.02 4.79
$\text{FL}^{2}\text{oRA}$-vision (Ours) 96.32 68.35 79.96 3.46 15.94 5.69 1.35 4.91 2.12 3.06 15.26 5.10
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 97.86 73.52 83.97 3.18 9.06 4.71 1.21 2.18 1.56 2.96 2.94 2.95
$\text{FL}^{2}\text{oRA}$-vision (Ours) 96.26 68.20 79.83 2.44 16.82 4.26 0.92 5.38 1.57 2.07 15.89 3.66
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 97.62 72.42 83.15 3.37 9.04 4.91 1.28 2.06 1.58 3.18 9.05 4.71
$\text{FL}^{2}\text{oRA}$-vision (Ours) 96.45 68.12 79.85 3.42 16.24 5.65 1.35 4.95 2.12 3.06 15.55 5.11
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 94.98 97.96 96.45 1.74 2.70 2.12 0.64 0.96 0.77 1.33 2.41 1.71
$\text{FL}^{2}\text{oRA}$-vision (Ours) 95.02 96.40 95.70 1.00 1.10 1.05 0.38 0.35 0.36 0.77 0.79 0.78
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 95.08 98.03 96.53 1.31 2.60 1.74 0.54 0.93 0.68 1.05 2.36 1.45
$\text{FL}^{2}\text{oRA}$-vision (Ours) 95.01 95.78 95.39 0.78 0.78 0.78 0.25 0.25 0.25 0.51 0.55 0.53
@@
$\text{FL}^{2}\text{oRA}$-text (Ours) 95.06 97.91 96.46 1.65 2.59 2.02 0.61 0.94 0.74 1.29 2.34 1.66
$\text{FL}^{2}\text{oRA}$-vision (Ours) 95.06 96.24 95.64 0.92 0.91 0.91 0.29 0.28 0.29 0.80 0.64 0.71

### v2

@@ 0.A.4 Calibration baselines
Eq. 10 computes an auxiliary calibration loss by taking, over $m$ samples in a minibatch, that penalizes the mismatch between observed mean correctness and mean confidence, thereby guiding the model to better align its confidence estimates with accuracy, without a strict constraint on each example. Note that during backpropagation, the gradients are backpropagated only through the confidence term $s_{i}$.
@@
Additional implementation details. In this paragraph, we elaborate on the implementation details, providing additional information beyond what is included in the main paper. In $\text{FL}^{2}\text{oRA}$, we only fine-tune the LoRA-adapted layers of the transformer blocks, while keeping all original weights frozen. Specifically, for both the image and text encoders, only the LoRA matrices $W_{A}$ and $W_{B}$ of the query, key, and value projections in the attention modules are trainable. All other parameters, including projection, MLP, and layer normalization weights, remain frozen to reduce memory footprint and improve training efficiency. We keep the rank of LoRA metrices to 2 and a dropout rate of 0.25 to regularize the input of the LoRA modules. Furthermore, we set the number of bins to $G=10$ for all the calibration metrics.
Appendix 0.B Additional empirical evaluation
In this section, we provide additional experimental analyses and a detailed breakdown of the results presented in the main paper.
Method FL setting (w/o calib) FL setting (w/ calib: DCA [dca]) FL setting (w/ calib: MDCA [mdca])
Brier Score $\downarrow$ NLL $\downarrow$ Brier Score $\downarrow$ NLL $\downarrow$ Brier Score $\downarrow$ NLL $\downarrow$
CIFAR-10
ZS-CLIP [radford2019language] 0.17 0.37 0.17 0.37 0.17 0.37
PromptFL [promptfl] 0.12 0.23 0.12 0.24 0.12 0.23
FedPGP [fedpgp] 0.13 0.28 0.13 0.28 0.14 0.30
FedOTP [fedotp] 0.09 0.20 0.09 0.20 0.09 0.20
FedPHA [fedpha] 0.10 0.20 0.10 0.20 0.10 0.20
$\text{FL}^{2}\text{oRA}$ (Ours) 0.05 0.09 0.05 0.09 0.05 0.09
CIFAR-100
ZS-CLIP [radford2019language] 0.48 1.40 0.48 1.40 0.48 1.40
PromptFL [promptfl] 0.36 0.93 0.36 0.92 0.36 0.92
FedPGP [fedpgp] 0.42 1.12 0.38 1.00 0.38 0.98
FedOTP [fedotp] 0.38 1.00 0.37 0.97 0.38 1.00
FedPHA [fedpha] 0.35 0.91 0.34 0.90 0.34 0.91
$\text{FL}^{2}\text{oRA}$ (Ours) 0.50 1.43 0.23 0.52 0.23 0.52
0.B.1 Additional calibration metrics
@@
In Table 9, we present the detailed Brier score and NLL calibration results of $\text{FL}^{2}\text{oRA}$ and the baseline methods. We observe that $\text{FL}^{2}\text{oRA}$ consistently achieve better calibration compared to the baselines. This is consistent with the findings in the main paper, where for all the calibration metrics, $\text{FL}^{2}\text{oRA}$ is better and/or competitive without needing extra regularization losses.
@@
0.B.2 Impact of client participation rate
@@
0.B.3 Detailed results of the main tables
In Table 10, 11 and 12, we provide the detailed results of individual datasets PACS, OfficeHome and VLCS respectively in the domain-generalization setting. The average results of them are reported in the Table. 4 of the main paper. Similarly in Table 13, 14, 15, 16 and 17, we provide the detailed results of Food101, DTD, Caltech101, Flowers102, and OxfordPets datasets, corresponding to Table 5 of the main paper, for the base-to-new generalization setting.